OpenStudy (anonymous):

A roof that encloses an attic is a square pyramid with a base edge length of 45 feet and a height of 5 feet. What is the volume of the attic in cubic feet and in cubic yards? (Hint: 27 cubic feet = 1 cubic yard.)

OpenStudy (jdoe0001):

 notice the base of the pyramid, is an SQUARE, thus a square pyramid \(\bf \textit{volume of a pyramid}=\cfrac{1}{3}\cdot \textit{squarish base}\cdot height\)

OpenStudy (anonymous):

so it would be 64.98??

OpenStudy (tkhunny):

Would what be 64.98? Please try to use complete notation. Volume: (1/3) * (45 ft*45 ft) * 5 ft = 5*15*45 ft^3 = 3,375 ft^3 See how that is labeled? All the calculations are right there, for everyone to see. No one has to guess what is meant. Your turn. Calculate the Cubic Yards.